I have a problem in the form of W=Cr^-2 and I need to solve for r. I am lost at this type of problem and could use some assistance.
Answer by checkley77(12844)

W=Cr^-2
W=C/r^2
Wr^2=C
r^2=C/W
r=sqrt(C/W)